# Formalizing Complex Mathematical Statements with LLMs: A Study on Mathematical Definitions

### v2

@@ 3.1.1 Error Analysis & Interventions
We observe that syntactic errors (SYN) exhibit the most variety, suggesting that GPT-4o may struggle to follow syntactic rules in Isabelle/HOL if not explicitly instructed. Type unification errors (TUF) suggest that GPT-4o my struggle with the exact usage of defined Isabelle items. To improve these issues, we investigate a Categorical Refinement (CR) method. CR involves designing specific additive instructions that constraint the behaviors leading to errors identified in the qualitative analysis.
Similarly, for syntactic errors (SYN), causes 1, 2, and 3 in Table 4 can be addressed with rule-based algorithms that refine formal codes at the symbolic level (Symbolic Refinement, SR).
Undefined errors (UDF), on the other hand, indicate that although GPT-4o can refer to external formal mathematical items, it remains unaware of the location or existence of relevant libraries. To alleviate UDF errors, we propose the process of Formal Definition Grounding (FDG) based on two methods: 1. Postprocessing (Post-FDG): explicitly augment preambles generated by LLMs with relevant libraries; 2. Prompting (Prompt-FDG): provide LLMs with grounded formal items and preambles in context to guide autoformalization.
3.2 Categorical Refinement
In order to better understand the refinement capabilities of GPT4-o, we investigate a set of error correction strategies: (i) Plain: provide LLMs with previously generated formal codes; (ii) Binary: additionally, provide LLMs with the correctness status of the formal code; (iii) Detailed: instead of just the binary correctness status, provide LLMs with the details of type, message, and line location of individual errors in the code.
In addition, to evaluate categorical refinement, we design specific instructions for each category of errors based on our qualitative analysis (Table 5). We report the error rate results of different refinement methods on GPT-4o in bar charts in Figure 2. All prompts used for categorical refinement along with additional empirical results are provided in Appendix.
@@ 3.2 Categorical Refinement
Categorical refinement demonstrates superiority in reducing error rates.
As shown in Figure 2(a), across all datasets, the refinement method that achieves the lowest overall error rate incorporates one of the instructions for categorical refinement, highlighting its superiority. However, when categorical refinement is applied without error details, such improvements do not occur. We hypothesize that this is because categorical instructions serve as constraints, making it more difficult for the LLM to follow them without more detailed error information for individual instances. Once such information is provided, the LLM receives sufficient information to adhere to the categorical refinement instructions.
(a) ZS on miniF2F-Test
(b) ZS on Def_Wiki-Test
@@
As shown in Figure 2(b), the method with the lowest SYN error rate on miniF2F-Test is plain refinement with SYN categorical refinement instructions, whereas on the other two datasets the best performing method is SYN categorical refinement with error details. In Figure 2(c), UDF categorical refinement with error details also leads to the lowest UDF error rate on all three datasets. Similarly in Figure 2(d), TUF categorical refinement with error details achieves the lowest TUF error rate on two out of the three datasets. These results collectively demonstrate the effectiveness of the design of categorical refinement. The only exceptional is TUF errors on the Def_ArXiv dataset, which again highlights the difficulty of fixing TUF errors.
@@ 3.3 Symbolic Refinement
The differences in error rates between our methods and direct testing results are shown as bar charts in Figure 3. Detailed numbers and additional results are provided in Appendix.
@@
In Figures 3(b) and 3(e), both applying SR alone and in combination with Post-FDG lead to a lower SYN error rate on Def_Wiki-Test. On Def_ArXiv, Figures 3(c) and 3(f) similarly shows that applying SR alone results in a reduction of SYN errors. These results suggest that SR is an effective approach for addressing SYN errors. On miniF2F-Test, however, SR does not influence the error rates. This is because SR is closely tied to specific error patterns in the dataset.
3.4 Formal Definition Grounding (FDG)
3.4.1 FDG via Post-Processing
For FDG, we first extracted external formal definitions of mathematical items and their sources from the Isabelle/HOL library. Then we filtered the extracted definitions to retain only those likely relevant to the autoformalization task on the datasets. Finally, for each individual instance in Def_Wiki and Def_ArXiv, we manually determined which formal definitions should be provided as contextual elements. For miniF2F, we simply selected the definitions of real and complex numbers as the relevant definitions.
Autoformalization performance can be underestimated without including contextual information. In Figure 3, without modifying the main body of the formalization, replacing the preambles with possible preambles via FDG (Post-FDG) directly leads to higher overall syntactic correctness. On miniF2F-Test, this setting only considers sources containing formal definitions of real and complex numbers, yet it increases overall syntactic correctness by more than 40%.
@@ 3.4.1 FDG via Post-Processing
In Figure 3(a) and Figure 3(d), Post-FDG leads to positive performance gains across all error categories. However, in Figures 3(b), 3(c), 3(e) and 3(f), while UDF error rates decrease, error rates in other categories can increase. A similar trend is observed when applying SR, where a reduction in SYN errors can coincide with increases in errors from the other two categories. This phenomenon suggests that because definition datasets are more complex, LLMs are more prone to generating entangled errors during the autoformalization process.
3.4.2 FDG via Prompting
@@
We designed two prompts to include external formal definitions for FDG: 1. Soft: allow the LLM some flexibility in whether to use in-context formal definitions for autoformalization; 2. Hard: explicitly instruct the LLM to use the in-context formal definitions if they are related. We tested these prompts on GPT-4o and Def_Wiki-Test to evaluate whether it can correctly refer to formalised items in context. The results are reported in Table 4.
Intuitively, LLMs should perform better when more relevant information is provided within the prompt. However, directly including grounded formal definitions does not positively impact the formalisation. This behaviour indicates that current state-of-the-art LLMs cannot effectively link to relevant in-context formal items for autoformalization.
4 Related Work
Autoformalization builds connections between natural language and formal languages Quan et al. (2024b, a). It also plays an important role in formal mathematical reasoning. For instance, proof autoformalization has been used as an intermediate step in automated theorem proving (Jiang et al., 2023; Tarrach et al., 2024). Deep learning models, such as transformers, have been applied to autoformalization. For example, Cunningham et al. (2022) developed a transformer-based model for autoformalizing of theorems along with their proofs in Coq. In recent years, with the increasing capabilities of LLMs, prompting-based methods have also demonstrated the ability to autoformalize mathematical statements in Isabelle (Wu et al., 2022; Zhang et al., 2024; Li et al., 2024) and Lean (Lu et al., 2024). Despite recent progress in autoformalization, few studies have analyzed this task from an error perspective. Our work aim to address this gap.
There are a few benchmarks providing informal-formal mathemtical statement pairs. miniF2F dataset (Zheng et al., 2022) provided 488 mathematical statement pairs from high school and undergraduate level to Olympiad problems in Lean, Metamath, and Isabelle. ProofNet (Azerbayev et al., 2023) benchmark contains 371 parallel formal theorem statements, natural language theorem statements, and natural language proofs in Lean. However, informal-Formal mathematical statement pairs are still scarce. Obtaining ground-truth formal codes requires specialists and there are many ways to formalize mathematical statements. In our work, we only provide the datasets in the natural language version and aim to develop methods without ground-truth formal codes.
5 Conclusion
This paper explored the challenges and advancements in autoformalization of complex mathematical statements. To this end, two datasets collecting real-world definitions in machine learning were introduced for systematic evaluation. By assessing autoformalization performance across three modern LLMs on newly introduced datasets, we identify key failure patterns including syntactic inconsistency, undefined references, and type mismatch. To address these, we proposed interventions such as Formal Definition Grounding and Categorical Refinement to enhance performance. Our results suggest that while modern LLMs exhibit potential in converting natural language mathematical definitions into formal representations, they still require improved guidance mechanisms and structured refinement techniques to enhance accuracy. Future research should focus on strengthening self-correction capabilities and integrating more robust contextual understanding into LLM-based formalization systems.
6 Limitations
Despite its contributions, this study has several limitations. First, the error analysis was conducted in Isabelle/HOL, and results may not directly generalize to other formal proof assistants such as Lean. Second, the definition datasets proposed, though diverse, are relatively small scale. Additionally, while the proposed refinements improve formalization performance, they do not fully eliminate semantic inconsistencies between natural language definitions and their formalized counterparts. More advanced methods are still needed to be developed.
Acknowledgements
This work was funded by the Swiss National Science Foundation (SNSF) project “NeuMath” (200021_204617), by the CRUK National Biomarker Centre, and supported by the Manchester Experimental Cancer Medicine Centre and the NIHR Manchester Biomedical Research Centre.
Appendix A Case Study for Formal Definition Grounding
Definition of Bradley–Terry model: Given a pair of items $i$ and $j$ drawn from some population, the Bradley–Terry model estimates the probability that the pairwise comparison turns out true, as $\Pr(i>j)=\frac{p_{i}}{p_{i}+p_{j}}$ where $p_{i}$ is a positive real-valued score assigned to individual $i$. Generated Formal Code: theory test imports Main begin definition bradley_terry :: "real \<Rightarrow> real \<Rightarrow> real" where "bradley_terry p_i p_j = p_i / (p_i + p_j)" end
Category Reasons
UDF 1. Items not defined. Formalization requires every mentioned item to be clearly defined in the local context or preambles. For one piece of generated formal code, GPT-4o could refer to items that are not defined in both sources.
The preamble in the generated formal code is “Main”. However, “Main” does not contain the formalization of “real”, making the formal code invalid. After applying Post-FDG, the preamble is updated to “HOL.Real”, and the formal code becomes valid. One might suggest creating a universal preamble that imports all source files from the library, applying this common preamble to solve such issues. However, this approach would not align with how a human expert would perform formalization. This failure to identify the correct preambles exposes limitations in the autoformalization capabilities of LLMs. Another issue, which is outside the scope of this paper but an important future direction, is that while Post-FDG can correct the formal code, the semantics of the generated code still do not fully match the original natural language version. For instance, the term “probability” does not appear in the formal code, and the phrase “$p_{i}$” is a positive real number” is omitted. The challenge of measuring semantic consistency between the generated formal code and its corresponding natural language version remains an open problem.
Appendix B Examples of Incorrect Formal Codes
In this section, we provide some examples of incorrect formal codes generated by GPT-4o to support our summarized reasons in Table 5. All examples of definitions are from Def_Wiki development set.
Example 1: Definition of Rademacher Complexity: Given a set $A\subseteq\mathbb{R}^{m}$, the Rademacher complexity of A is defined as follows: $\operatorname{Rad}(A):=\frac{1}{m}\mathbb{E}_{\sigma}\left[\sup_{a\in A}\sum_{i=1}^{m}\sigma_{i}a_{i}\right]$ where $\sigma_{1},\sigma_{2},\dots,\sigma_{m}$ are independent random variables drawn from the Rademacher distribution (i.e. $\Pr(\sigma_{i}=+1)=\Pr(\sigma_{i}=-1)=1/2$ for $i=1,2,\dots,m$), and $a=(a_{1},\dots,a_{m})$. Generated Formal Code: theory test imports Main begin definition rademacher_distribution :: "nat \<Rightarrow> real pmf" where "rademacher_distribution m = pmf_of_set {x. x = 1 \<or> x = -1}" definition rademacher_complexity :: "(real vector set) \<Rightarrow> nat \<Rightarrow> real" where "rademacher_complexity A m = (1 / real m) * (expectation (pmf_of_set {x. x = 1 \<or> x = -1}) (\<lambda>\<sigma>. Sup {sum (\<lambda>i. \<sigma i * a i) {1..m} | a. a \<in> A}))" end
Example 2: Definition of Algorithm Selection: Given a portfolio $\mathcal{P}$ of algorithms $\mathcal{A}\in\mathcal{P}$, a set of instances $i\in\mathcal{I}$ and a cost metric $m:\mathcal{P}\times\mathcal{I}\to\mathbb{R}$, the algorithm selection problem consists of finding a mapping $s:\mathcal{I}\to\mathcal{P}$ from instances $\mathcal{I}$ to algorithms $\mathcal{P}$ such that the cost $\sum_{i\in\mathcal{I}}m(s(i),i)$ across all instances is optimized. Generated Formal Code: theory test imports Main begin type_synonym ${}^{\prime}\!$a instance = ${}^{\prime}\!$a type_synonym ${}^{\prime}\!$a algorithm = ${}^{\prime}\!$a type_synonym cost_metric = "${}^{\prime}\!$a algorithm \<Rightarrow> ${}^{\prime}\!$a instance \<Rightarrow> real" definition algorithm_selection :: "${}^{\prime}\!$a algorithm set \<Rightarrow> ${}^{\prime}\!$a instance set \<Rightarrow> cost_metric \<Rightarrow> (${}^{\prime}\!$a instance \<Rightarrow> ${}^{\prime}\!$a algorithm) \<Rightarrow> real" where "algorithm_selection \<mathcal{P} \<mathcal{I} m s = (\<Sum>i \<in> \<mathcal{I}. m (s i) i)" end
Example 3: Definition of Set-Family Growth Function: Let $H$ be a set family (a set of sets) and $C$ a set. Their intersection is defined as the following set-family: $H\cap C:=\{h\cap C\mid h\in H\}$. The intersection-size (also called the index) of $H$ with respect to $C$ is $|H\cap C|$. The growth function measures the size of $H\cap C$ as a function of $|C|$. Formally: $\operatorname{Growth}(H,m):=\max_{C:|C|=m}|H\cap C|$. Generated Formal Code: theory test imports Main begin definition intersection_set_family :: "${}^{\prime}\!$a set set \<Rightarrow> ${}^{\prime}\!$a set \<Rightarrow> ${}^{\prime}\!$a set set" (infixl "\<inter>\<^sub>H" 70) where "H \<inter>\<^sub>H C = {h \<inter> C | h. h \<in> H}" definition intersection_size :: "${}^{\prime}\!$a set set \<Rightarrow> ${}^{\prime}\!$a set \<Rightarrow> nat" where "intersection_size H C = card ( H \<inter>\<^sub>H C)" definition growth_function :: "${}^{\prime}\!$a set set \<Rightarrow> nat \<Rightarrow> nat" where "growth_function H m = (SUP C. card C = m \<and> finite C \<and> (\<forall>h \<in> H. finite h) \<and> finite H \<longrightarrow> card (H \<inter>\<^sub>H C))" end
Appendix C Prompts and Detailed Results
The prompts used for the estimation of dataset statistics are provided in Table 6. The instructions used in the prompts of experiments are provided in Table 7. Detailed numbers of autoformalization results on miniF2F test set, Def_Wiki test set and Def_ArXiv are provided in Table 8, 9, 10, respectively. Symbolic refinement results and Post-FDG results on Def_Wiki test set are provided in Table 11 and Table 12, respectively.

### v3

@@ 3.2 Categorical Refinement
In order to better understand the refinement capabilities of GPT-4o, we investigate a set of error correction strategies: (i) Plain: provide LLMs with previously generated formal codes; (ii) Binary: additionally, provide LLMs with the correctness status of the formal code; (iii) Detailed: instead of just the binary correctness status, provide LLMs with the details of type, message, and line location of individual errors in the code.
In addition, to evaluate categorical refinement, we design specific instructions for each category of errors based on our qualitative analysis (Table 3). We report the error rate results of different refinement methods on GPT-4o in bar charts in Figure 2. All prompts used for categorical refinement and additional empirical results are provided in Appendix.
@@
Categorical refinement reduces error rates.
As shown in Figure 2(a), across all datasets, the refinement method that achieves the lowest overall error rate incorporates one of the instructions for categorical refinement, highlighting the efficacy of this mechanism. However, when categorical refinement is applied without error details, such improvements do not occur. We hypothesize that this is because categorical instructions serve as constraints, making it more difficult for the target LLM to follow them without more detailed error information for individual instances. Once such information is provided, the LLM receives sufficient information to adhere to the categorical refinement instructions.
As shown in Figure 2(b), the method with the lowest SYN error rate on miniF2F-Test is plain refinement with SYN categorical refinement instructions, whereas on the other two datasets the best performing method is SYN categorical refinement with error details. In Figure 2(c), UDF categorical refinement with error details also leads to the lowest UDF error rate on all three datasets. Similarly in Figure 2(d), TUF categorical refinement with error details achieves the lowest TUF error rate on two out of the three datasets. These results collectively demonstrate the effectiveness of the categorical refinement as a control mechanism for autoformalization. The only exception is TUF errors within the Def_ArXiv dataset, which again highlights the difficulty of fixing TUF errors.
(a) ZS on miniF2F
(b) ZS on Def_Wiki
@@ 3.3 Symbolic Refinement
The differences in error rates between our methods and direct testing of unmodified autoformalized code are illustrated as bar charts in Figure 3.
@@
In Figures 3(b) and 3(e), both applying symbolic refinement (SR) alone and in combination with Post-FDG lead to a lower SYN error rate on Def_Wiki-Test. On Def_ArXiv, Figures 3(c) and 3(f) similarly shows that applying SR alone results in a reduction of SYN errors. These results suggest that SR is an effective approach for addressing SYN errors. On miniF2F-Test, however, SR does not impact the error rates because SR is closely tied to specific error patterns in the dataset.
3.4 Post-FDG
For implementing FDG, we first extracted external formal definitions of mathematical items and their sources from the Isabelle/HOL library. Then we filtered the extracted definitions to retain only those likely relevant to the autoformalization task on the datasets. Finally, for each individual instance in Def_Wiki and Def_ArXiv, we manually determined which formal definitions should be provided as contextual elements. For miniF2F, we simply selected the definitions of real and complex numbers as the relevant definitions. Post-FDG (FDG via postprocessing) explicitly augments the preambles generated by LLMs with the sources of relevant formal definitions in formal libraries.
Autoformalization performance can be underestimated without including contextual information.
In Figure 3, without modifying the main body of the formalization, replacing the preambles with possible preambles via Post-FDG directly leads to higher overall syntactic correctness. On miniF2F-Test, this setting only considers sources containing formal definitions of real and complex numbers, yet it increases overall syntactic correctness by more than 40%.
@@ 3.4 Post-FDG
In Figure 3(a) and Figure 3(d), Post-FDG leads to positive performance gains across all error categories. However, in Figures 3(b), 3(c), 3(e) and 3(f), while UDF error rates decrease, error rates in other categories can increase. A similar trend is observed when applying SR, where a reduction in SYN errors can coincide with increases in errors from the other two categories. This phenomenon suggests that because definition datasets are more complex, LLMs are more prone to generating errors from different categories in one code block during the autoformalization process.
3.5 Generalizability to Lean4
Prompt Strategy miniF2F Def_Wiki Def_Arxiv
ZS 13.93 5.36 0.00
(ZS)+D 15.98 7.14 6.67
(ZS)+D+CR-SYN 16.80 8.93 6.67
(ZS)+D+CR-UDF 13.52 10.71 6.67
(ZS)+D+CR-TUF 15.57 10.71 6.67
Table 4: GPT-4o pass rates with Lean4. ((ZS)+D): detailed refinement on (zero-shot) formal codes.
We further explore the generalizability of our methods to formal languages beyond Isabelle/HOL. We select Lean4 (de Moura et al., 2015) as a representative target due to its increasing popularity and widespread use. The Categorical Refinement (CR) method can be applied to Lean4 with only minor prompt modifications. In contrast, exploring the generalizability of Symbolic Refinement and Formal Definition Grounding requires system-specific designs, which fall outside the scope of this paper. We investigate the generalizability of CR on Lean4 and report the pass rates across three datasets using various strategies in Table 4.
Definitions present greater complexity for autoformalization in Lean4.
The pass rates of GPT-4o on definitions are consistently lower than those on miniF2F. For example, on Def_Arxiv, GPT-4o fails to correctly formalize any of the 30 definitions. Moreover, we observe that providing error details to revise the output improves pass rates, aligning with our observations in Isabelle/HOL.
Categorical Refinement generalizes to Lean4.
On the miniF2F dataset, CR-SYN achieves the highest pass rate. Similarly, on the Def_Wiki dataset, CR-UDF and CR-TUF yield the best results. These findings demonstrate the effectiveness of the proposed CR in other formal languages. Notably, since the Lean4 assistant does not provide explicit categories as Isabelle/HOL, performance differences across CR categories may indirectly reflect distinct types of errors encountered during autoformalization.
4 Related Work
Autoformalization allows for a systematic connection between material and formal inferences Quan et al. (2024a, b), also enabling the universalization of formal mathematical reasoning. For instance, proof autoformalization has been used as an intermediate step in automated theorem proving (Jiang et al., 2023; Tarrach et al., 2024). Deep learning models, such as transformers, have been applied to autoformalization in Coq (Cunningham et al., 2022). In recent years, with the increasing capabilities of LLMs, prompting-based methods have also demonstrated the ability to autoformalize mathematical statements in Isabelle (Wu et al., 2022; Zhang et al., 2024; Li et al., 2024) and Lean (Yang et al., 2023; Lu et al., 2024; Liu et al., 2025a). Despite recent progress in autoformalization with LLMs, few studies have analyzed this task from an error perspective. Our work takes a step in this direction.
There are a few benchmarks that provide informal–formal mathematical statement pairs. MiniF2F (Zheng et al., 2022) and ProofNet benchmark (Azerbayev et al., 2023) include samples paired with ground-truth formal statements ranging from high-school and undergraduate problems to Olympiad-level problems. However, such informal–formal pairs remain scarce. A growing trend is the development of data generation pipelines for constructing large-scale parallel corpora to finetune LLMs for autoformalization (Jiang et al., 2024; Liu et al., 2025b). While useful, these benchmarks still focus on exercise-style mathematical problems, which do not fully reflect real-world scenarios. In contrast, our definition datasets emphasize real-world statements.
5 Conclusion
This paper explored the challenges and advancements in autoformalization of complex mathematical statements. To this end, two datasets collecting real-world definitions in machine learning were introduced for systematic evaluation. By assessing autoformalization performance across three modern LLMs on newly introduced datasets, we identify key failure patterns including syntactic inconsistency, undefined references, and type mismatch. To address these, we proposed interventions such as Categorical Refinement and Formal Definition Grounding to enhance performance. Our results suggest that while modern LLMs exhibit potential in converting natural mathematical definitions into formal representations, they still require improved guidance mechanisms and structured refinement techniques to enhance accuracy. Future research could focus on improving self-correction capabilities and integrating more robust contextual understanding into LLM-based formalization systems.
6 Limitations
Despite its contributions, this study has several limitations. First, the error analysis was conducted in Isabelle/HOL, and some results may not directly generalize to other formal proof assistants such as Lean. Second, the definition datasets proposed, though diverse, are relatively small scale. Additionally, while the proposed refinements improve formalization performance, they do not fully eliminate semantic inconsistencies between natural language definitions and their formalized counterparts. More advanced methods are still needed to be developed.
Acknowledgements
This work was funded in part by the Swiss National Science Foundation (SNSF/FAPESP) “RATIONAL” project.
Appendix A Detailed Information about the Dataset Creation
We obtain mathematical definitions in the machine learning domain from two sources: Wikipedia (Def_Wiki) and Arxiv Papers (Def_ArXiv). For Def_Wiki, definitions are collected from pages under the Machine Learning category [^2] https://en.wikipedia.org/wiki/Category:Machine_learning and its sub-categories. We manually browsed each page, identified well-defined definitions (i.e., formal descriptions with mathematical symbols), and converted the chosen definitions into LaTeX format. In total, we obtained 56 qualified natural language definitions in LaTeX and divided them into development and test sets, containing 10 and 46 samples, respectively. For Def_ArXiv, we used the advanced search tool on ArXiv’s website, filtering for papers in the subcategories cs.LG and stat.ML, with comments including "ICML." We restricted the search to papers published in 2019, 2020, and 2021 and manually reviewed the first 25 papers from each year. We shortlisted papers that were accepted to the ICML conference and contained formally described definitions with mathematical symbols to ensure reliability. We then filtered out definitions that were less straightforward or formal in their expressions, extracted the LaTeX versions, and ultimately obtained 30 definitions from 7 papers. We provide dataset statistics in Table 5.
Property miniF2F-Test Def_Wiki Def_ArXiv
No. Samples 244 56 30
No. Tokens 70.25 (47.70) 200.18 (112.98) 164.40 (71.47)
No. Objects 4.76 (1.68) 7.63 (2.71) 7.10 (2.64)
No. Formulae 2.71 (1.74) 2.84 (2.05) 3.17 (1.97)
Appendix B Case Study for Formal Definition Grounding
Definition of Bradley–Terry model:
Given a pair of items $i$ and $j$ drawn from some population, the Bradley–Terry model estimates the probability that the pairwise comparison turns out true, as
$\Pr(i>j)=\frac{p_{i}}{p_{i}+p_{j}}$
where $p_{i}$ is a positive real-valued score assigned to individual $i$.
theory test
imports Main
begin
definition bradley_terry :: "real \<Rightarrow> real
\<Rightarrow> real" where "bradley_terry p_i p_j =
p_i / (p_i + p_j)"
end
The preamble in the generated formal code is “Main”. However, “Main” does not contain the formalization of “real”, making the formal code invalid. After applying Post-FDG, the preamble is updated to “HOL.Real”, and the formal code becomes valid. One might suggest creating a universal preamble that imports all source files from the library, applying this common preamble to solve such issues. However, this approach would not align with how a human expert would perform formalization. This failure to identify the correct preambles exposes limitations in the autoformalization capabilities of LLMs. Another issue, which is outside the scope of this paper but an important future direction, is that while Post-FDG can correct the formal code, the semantics of the generated code still do not fully match the original natural language version. For instance, the term “probability” does not appear in the formal code, and the phrase $p_{i}$” is a positive real number” is omitted.
We acknowledge the importance of quantitatively evaluating semantic consistency in autoformalization. In this paper we mainly quantitatively evaluate the syntactic aspects of formalizations because syntactic correctness can be systematically and fully validated through the theorem prover. However, the full evaluation of semantic consistency still presents technical challenges and remains an open research question. Additionally, it is important to notice that semantic validity is not completely disentangled from syntactic checks. Some important semantic aspects, in fact, are explicitly or implicitly covered within our evaluation. It can be noted, for example, that some of the errors in Table 5 also cover semantic aspects. For instance, a model that is unaware of undefined items will also likely lack a complete semantic understanding of the crucial elements required for correct autoformalization. We believe that coming up with a reliable quantitative method to systematically assess semantic alignment is an important research question to explore in future work, which is out of scope for the current paper.
Appendix C Examples of Incorrect Formal Codes
We provide some examples of incorrect formal codes generated by GPT-4o to support our summarized reasons in Table 3. All examples of definitions are from Def_Wiki development set.
Example 1:
Definition of Rademacher Complexity:
Given a set $A\subseteq\mathbb{R}^{m}$, the Rademacher complexity of A is defined as follows:
$\operatorname{Rad}(A):=\frac{1}{m}\mathbb{E}_{\sigma}\left[\sup_{a\in A}\sum_{i=1}^{m}\sigma_{i}a_{i}\right]$
where $\sigma_{1},\sigma_{2},\dots,\sigma_{m}$ are independent random variables drawn from the Rademacher distribution (i.e. $\Pr(\sigma_{i}=+1)=\Pr(\sigma_{i}=-1)=1/2$ for $i=1,2,\dots,m$), and $a=(a_{1},\dots,a_{m})$.
theory test
imports Main
begin
definition rademacher_distribution ::
"nat \<Rightarrow> real pmf" where
"rademacher_distribution m =
pmf_of_set {x. x = 1 \<or> x = -1}"
definition rademacher_complexity ::
"(real vector set) \<Rightarrow> nat \<Rightarrow> real"
where "rademacher_complexity A m =
(1 / real m) * (expectation
(pmf_of_set {x. x = 1 \<or> x = -1})
(\<lambda>\<sigma>. Sup {sum (\<lambda>i. \<sigma i * a i)
{1..m} | a. a \<in> A}))"
end
Example 2:
Definition of Algorithm Selection:
Given a portfolio $\mathcal{P}$ of algorithms $\mathcal{A}\in\mathcal{P}$, a set of instances $i\in\mathcal{I}$ and a cost metric $m:\mathcal{P}\times\mathcal{I}\to\mathbb{R}$, the algorithm selection problem consists of finding a mapping $s:\mathcal{I}\to\mathcal{P}$ from instances $\mathcal{I}$ to algorithms $\mathcal{P}$ such that the cost $\sum_{i\in\mathcal{I}}m(s(i),i)$ across all instances is optimized.
theory test
imports Main
begin
[…]
definition algorithm_selection ::
[…]
"algorithm_selection \<mathcal{P}
\<mathcal{I} m s = (\<Sum>i \<in>
\<mathcal{I}. m (s i) i)"
end
Example 3:
Definition of Set-Family Growth Function:
Let $H$ be a set family (a set of sets) and $C$ a set. Their intersection is defined as the following set-family: $H\cap C:=\{h\cap C\mid h\in H\}$. The intersection-size (also called the index) of $H$ with respect to $C$ is $|H\cap C|$. The growth function measures the size of $H\cap C$ as a function of $|C|$. Formally: $\operatorname{Growth}(H,m):=\max_{C:|C|=m}|H\cap C|$.
theory test
imports Main
begin
definition intersection_set_family ::
[…]
(infixl "\<inter>\<^sub>H" 70) where
"H \<inter>\<^sub>H C = {h \<inter> C | h. h \<in> H}"
definition intersection_size ::
[…]
"intersection_size H C = card (
H \<inter>\<^sub>H C)"
[…]
nat \<Rightarrow> nat" where
"growth_function H m = (SUP C. card C =
m \<and> finite C \<and> (\<forall>h \<in> H. finite h)
\<and> finite H \<longrightarrow> card (H \<inter>\<^sub>H C))"
end
Appendix D Formal Definition Grounding via Prompting
We further explore formal definition grounding with prompting (Prompt-FDG): provide LLMs with grounded formal items and preambles in context to guide autoformalization. We designed two prompts to include external formal definitions for FDG: 1. Soft: allow the LLM some flexibility in whether to use in-context formal definitions for autoformalization; 2. Hard: explicitly instruct the LLM to use the in-context formal definitions if they are related. We tested these prompts on GPT-4o and Def_Wiki-Test to evaluate whether it can correctly refer to formalised items in context. The results are reported in Table 6.
Intuitively, LLMs should perform better when more relevant information is provided within the prompt. However, directly including grounded formal definitions does not positively impact the formalization. This behaviour indicates that current state-of-the-art LLMs cannot effectively link to relevant in-context formal items for autoformalization. How to successfully leverage in-context formal definitions for autoformalization with LLMs remains an important open research question.
Appendix E Prompts and Additional Results
The prompts used for the estimation of dataset statistics are provided in Table 7. The instructions used in the prompts of experiments are provided in Table 8. Detailed numbers of autoformalization results on miniF2F test set, Def_Wiki test set and Def_ArXiv are provided in Table 9, 10, 11, respectively. Symbolic refinement results and Post-FDG results on Def_Wiki test set are provided in Table 12 and Table 13, respectively.